OpenStudy (anonymous):

Rationalize the denominator and simply: square root of 5/2?

OpenStudy (amistre64):

\[\frac{\sqrt{5}}{2} ??\]

OpenStudy (amistre64):

\[\sqrt{\frac{5}{2}}??\]

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

yes to the second one I assume ;)

OpenStudy (anonymous):

It's the first one

OpenStudy (amistre64):

\[\sqrt{\frac{5}{2}} =\frac{\sqrt{5}}{\sqrt{2}}\]

OpenStudy (amistre64):

the first one is already rationalized and simplified..... it makes no sense to "do" it lol

OpenStudy (anonymous):

amistre64 is correct there is no way to go further on this problem

OpenStudy (amistre64):

\[\frac{\sqrt{5}}{\sqrt{2}} * \frac{\sqrt{2}}{\sqrt{2}} = \frac{\sqrt{10}}{2}\]

OpenStudy (anonymous):

How did you draw a fration with denominator on the next line? HAHAH.

OpenStudy (amistre64):

I know right? lol use fract{top}{bottom} in the equation editor..

OpenStudy (anonymous):

\[2\sqrt{10}/5\]

OpenStudy (anonymous):

\[\sqrt{5} / \sqrt{2}\] ooh thanks. I always get this

OpenStudy (amistre64):

\[\frac{numerator}{denominator}\]

OpenStudy (amistre64):

"frac" not "fract" ;)

OpenStudy (anonymous):

\[\frac{5}{2}\] LOL. It's working now. THANK YOU

OpenStudy (amistre64):

youre welcome :)